Determine all the first and second order partial derivatives for the function:

u(x,t)=Ce^(1-n²π²)t sin (nπx)


1
Expert's answer
2021-05-13T18:16:06-0400

Let us determine all the first order partial derivatives for the function

"u(x,t)=Ce^{(1-n²π²)t} \sin (nπx):"


"\frac{{\partial}u(x,t)}{\partial x}=Cn\pi e^{(1-n²π²)t} \cos (nπx), \ \ 
\frac{{\partial}u(x,t)}{\partial t}=C(1-n²π²)e^{(1-n²π²)t} \sin (nπx)."


Let us determine all the second order partial derivatives for the function

"u(x,t)=Ce^{(1-n²π²)t} \sin (nπx):"


"\frac{{\partial}u^2(x,t)}{\partial x^2}=-Cn^2\pi^2 e^{(1-n²π²)t} \sin (nπx), \ \ 
\frac{{\partial}u^2(x,t)}{\partial t^2}=C(1-n²π²)^2e^{(1-n²π²)t} \sin (nπx),"


"\frac{{\partial}u^2(x,t)}{\partial x \partial t}=Cn\pi(1-n²π²) e^{(1-n²π²)t} \cos (nπx), \ \ 
\frac{{\partial}u^2(x,t)}{\partial t \partial x}=Cn\pi (1-n²π²)e^{(1-n²π²)t} \cos (nπx)."